OpenStudy (anonymous):

Verify the identity. Show your work. cos(α - β) - cos(α + β) = 2 sin α sin β my answer>cos=o

OpenStudy (anonymous):

Teacher:Use the sum and difference formulas for this me:wtf you mean

OpenStudy (anonymous):

@misty1212

OpenStudy (sohailiftikhar):

cos(a-b)=cos(a)cos(b)+sin(a)sin(b) cos(a+b)=cos(a)cos(b)-sin(a)sin(b) these are the formulas

OpenStudy (anonymous):

woow ok so do i use both of them or what

OpenStudy (sohailiftikhar):

you have both terms cos(a-b) and cos(a+b) so use both

OpenStudy (sohailiftikhar):

in this problem

OpenStudy (anonymous):

ok so this is a practice question imma gonna be graded on so can you help me plz

OpenStudy (sohailiftikhar):

ok

OpenStudy (anonymous):

2cos(A)cos(B)=2cos(A)cos(B) ok so i got this instead is this right

OpenStudy (anonymous):

@misty

OpenStudy (sohailiftikhar):

no it's not right

OpenStudy (sohailiftikhar):

ok i'm going to solve it

OpenStudy (sohailiftikhar):

\[\cos(\alpha-\beta)-\cos(\alpha+\beta)=(\cos \alpha \cos \beta +\sin \alpha \sin \beta)-(\cos \alpha \cos \beta -\sin \alpha \sin \beta )\]

OpenStudy (sohailiftikhar):

now \[\cos \alpha \cos \beta +\sin \alpha \sin \beta -\cos \alpha \cos \beta +\sin \alpha \sin \beta =2 \sin \alpha \sin \beta \]